OpenStudy (anonymous):

1. Suppose you are at the gas station filling your tank with gas. The function C(g) represents the cost C of filling up the gas tank with g gallons. Given the equation: C(g) = 2.75(g) a. a. Find C(10) Show your work. b. b. Find C(15) Show your work. c. Will the domain for this function be all real numbers? Why or why not? 2. Examine the rise in gasoline prices from 2006 to 2009. The average price of regular unleaded gasoline in January 2006 was $2.31 and in January 2009 the average price of regular unleaded gasoline was $3.42. a.

OpenStudy (saifoo.khan):

a. C(g) = 2.75(g) C(10) = 2.75(10) C(10) = $27.5

OpenStudy (anonymous):

saifoo i need u

OpenStudy (saifoo.khan):

Hi Wlbrt. Sure, will be there in a sec.

OpenStudy (anonymous):

k

OpenStudy (saifoo.khan):

b. C(15) = 2.75(g) C(15) = 2.75(15) C(15) = $41.25

OpenStudy (saifoo.khan):

c. Yes it must be real numbers, this is b/c The gallons of petrol can't be a negative number